SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Nov.15, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dan Walker.

Operations

  • waiting to have colliding beam data.

EMT

  • EMT prescale is currently set to 1 till OPR release is ready to perform monitoring. Dan already has the OPR code ready, and it will be in the same release that contains the turn-on curve as well.
  • It's reminded to run the configuration tools on production area.

DCT/DCZ

  • The new TSF firmware is still under testing. Rainer still suggested Marc to re-route the old firmware to make sure that the problem wasn't due to the tools.

Online

  • Fang updated the DCZ online library against the latest online release yesterday. However, this version still doesn't save input TC for rare problems that would be useful for debugging for the future. Fang will test the code with a broken version of the new firmware so that we can make sure that L3 doesn't crash on that kind of damaged events. This can be tested at the test stand.

L3

  • Next week, Rainer will talk more about the DCH structure code. The timing plots look different when running Moose twice. He will fix this.
  • Rainer had installed new versions of L3PCASERVER. See HN posted.
  • Rainer is getting Arik on track for L3 tracking. Arik might have something to report next week.
  • Al is working on the trickle monitoring under the L3T pages in JAS. Some issues were discussed:
    • Need to check when the last reference plots were made and what conditions they belonged. They might need to be updated if necessary.
    • The xml files generated by the automatic script are removed and abandoned from the L3TMontool package. Al showed us a version of the L3T JAS page after the clean-up with more descriptions of the trickle injection plots. He will be checking in this version soon. After this, improvements on the L3T page can be made possible.
  • Prafulla is looking into OPR MON to fix the efficiency plot.

Data Quality and Monitoring

  • Kris and Rahmat are checking Run 4 reprocessed runs and found some problematic runs as showed here. Also see Kris' HN posting after the meeting.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.